OpenStudy (anonymous):

A map has a scale 1 in 38 ft. use the given map distance to find the actual map with 5.5 in. 2.25in 1.75 in. (inch)

OpenStudy (anonymous):

5.5 2.25,1.75 inches it different answer

OpenStudy (gw2011):

If the scale is 1 inch = 38 feet, then multiply each amount by 38. 5.5 in=(5.5)(38)=209 ft 2.25 in=(2.25)(38)=85.5 ft 1.75 in=(1.75)(38)=66.5 ft

OpenStudy (anonymous):

do you know about the another question 2.2/3.3=a-2/a-1

OpenStudy (gw2011):

2.2/3.3=(a-2)/(a-1) Cross multiply to get: (3.3)(a-2)=(2.2)(a-1) 3.3a-6.6=2.2a-2.2 3.3a-2.2a=6.6-2.2 1.1a=4.4 a=4.4/1.1 a=4

OpenStudy (anonymous):

thank you so much

OpenStudy (gw2011):

You're welcome